A rail road of length 20m is heated through 30°c to 100°c. calculate the gap between the rails

1
Expert's answer
2022-01-25T06:50:04-0500

Given:

"l_0=20\:\rm m"

"\Delta T=70^\circ"

"\alpha=12*10^{-6}\rm\:K^{-1}"


The change in the linear dimension of a rail due thermal expansion

"\Delta l=l_0\alpha\Delta T"

Hence, the gap between rails

"\Delta l=20*12*10^{-6}*70=0.0168\:\rm m=1.68\: cm"